OpenStudy (anonymous):

how would i find the radius and diameter of a circle whose circumfrence is 30 meters

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[circumference =2\pi r\]\[diameter =2r\]

myininaya (myininaya):

30=2pi(radius) => radius=30/(2pi)=15/pi 2(radius)=diameter =>diameter=2(15/pi)=30/pi